Consider the probability that fewer than 56 out of 104 people have not been in a car accident. Assume the probability that a given person has not been in a car accident is 21%. Specify whether the normal curve can be used as an approximation to the binomial probability by verifying the necessary conditions. ANSWER IS YES OR NO
Answer #1

np = 104*0.21 = 21.84

nq = 104*(1 - 0.21) = 82.16

Since np and nq both are greater than 5, we can use normal approximation.

The answer is Yes